In a recent survey of 205 never-married single women, 78% thought they would eventually find, and marry, their perfect mate. A 95% Confidence Interval for this proportion would be .78 plus or minus

1.645

.0476

.0567

.0157
Answer #1

Margin of error
(d - 1)*d *Z = 2
10.78 *(1-0.78) = 1.96 +1 205
=0.0567

A 95% Confidence Interval for this proportion would be .78 plus or minus 0.0567